# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v2

@@ 8.2 Scene and Demonstration Generation
To create simulation scenes for manipulation tasks, two methods can be used. Real-to-Sim method converts real scenes to simulation. Task-specific automation method generates task-specific simulation scenes automatically. Real-to-Sim method can accurately mimic the real world, but it limits the diversity of scenes. The task-specific automation method can create more diverse scenarios. This approach increases the variety of collected demonstrations.
The Real-to-Sim method needs to reconstruct the real scene (in Sec. 6) and to determine the positions of objects. 3D meshes of objects can be made from point cloud data or from RGB images using a generative model (Chen et al. (2024b)). But this Real-to-Sim method for generating object 3D meshes still fails to capture the object’s material and physical properties accurately such as liquid object.
The task-specific automation method requires specific assets tailored to the task. These assets must be imported and sized appropriately, with the scene also being configured to suit the task. RoboGen (Wang et al. (2023b)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023a)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library.
@@
To collect demonstrations in simulations, different approaches can be used based on task complexity. For simple tasks, like a two-finger gripper picking up a cube, a hard-coding method (Wang et al. (2022b)) can be used. However, for more complex tasks, remote teleoperation (Chen et al. (2024a)) or skill library (Ha et al. (2023)) should be employed. Building skill library can be done using reinforcement learning or gradient optimization methods. RoboGen (Wang et al. (2023b)) shows that gradient-based trajectory optimization is better for fine-grained manipulation tasks with soft bodies, like shaping dough into a specific form. On the other hand, reinforcement learning and evolutionary strategies are more effective for contact-rich tasks and continuous interactions with other components in the scene.
@@ 8.3 Sim-to-Real Gap Solutions
The methods discussed above assume that the target domain remains unchanged. However, many physical parameters of the same robot can change significantly. Factors like temperature, humidity, positioning, and wear and tear over time can all affect these parameters. This makes it harder to bridge the sim-to-real gap. To address this issue, DORA (Zhang et al. (2024b)) uses an information bottleneck principle. It aims to maximize the mutual information between the dynamics encoding and environmental data. At the same time, it minimizes the mutual information between the dynamics encoding and the behavior policy actions. Transic (Jiang et al. (2024)) proposes a data-driven approach that enables successful sim-to-real transfer using a human-in-the-loop framework.
@@ 8.4 Dataset Augmentation
Current dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s visual or pose (Mandlekar et al. (2023); Mandi et al. (2022)). GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
@@
Figure 8: Foundation Models for Manipulation Data Generation. Current mainstream simulators include Pybullet, MuJoCo, CoppeliaSim, NVIDIA Omniverse, and Unity. Meanwhile, learning-based generative models used as simulators have shown potential. Simulation environment generation can be classified into Real-to-Sim and Task-specific automation methods. In the Real-to-Sim method, assuming the object’s position is known, the main challenge lies in constructing the object’s 3D mesh. This can be achieved through scanning technique or by using VGM to generate the 3D mesh directly from RGB image (Chen et al. (2024b)). In the Task-specific automation method, LLM can output scene descriptions or scene code based on task instruction. When the output is a scene description, VGM generates the objects and arranges them according to the description. Meanwhile, the generated scene need to be evaluated by VLM. (Wang et al. (2023b)). When the output is scene code, it directly generates the corresponding scene (Wang et al. (2023a)). However, this requires substantial prior knowledge of scene code within the task library. There are three methods for generating demonstrations in a scene: Hard-code, Teleoperation, and Skill Library. When building skill library, gradient optimization is effective in training skill for deformable tasks and reinforcement learning works better for contact-rich tasks (Wang et al. (2023b)). Solutions for the Sim-to-Real gap include System Identification, Domain Randomization, and Transfer Learning. For data augmentation, VFM is used to segment images first, and then VGM renders the object’s texture on the masked image.
@@ 9.1 What is the framework for general manipulation?
Based on the definition and considerations of safety and accuracy at the L0 level, this paper proposes a framework for a general manipulation capability. Given that the scenarios are static, the framework is designed in a modular, sequential manner. To facilitate module migration, it is preferable for each module to be plug-and-play. Given the current reliance on human-in-the-loop mechanisms in autonomous driving and medical robotics to ensure safety, this framework aims for human-robot interaction through corrective instruction to ensure the safety of manipulation actions. The corrective action can be collected into the dataset and then improve old skills through offline training.
@@
During robot execution, continuous human supervision is not always feasible. Hence, integrating real-time monitoring through parallel surveillance videos during robot execution could enhance safety. The framework in this paper does not explicitly denote this parallel safety monitoring module, as it resembles the post-conditions detection module. The post-conditions detection module analyzes the robot’s execution video to identify reasons for task failure, facilitating post-hoc correction to ensure task success. If the algorithm’s task execution safety is 80%, and the monitoring module predicts safety at 80% as well, the probability of risky movements reduces to 4%. Of course, for household robots, ensuring an over 99% safety rate is imperative. Initially, cloud-based monitoring of multiple robots by a single operator, with human intervention to correct erroneous behaviors, appears to be the best approach. This strategy not only reduces labor requirements but also ensures safety. Later, by gathering extensive data to improve model accuracy. This framework still has shortcomings, such as the sequential execution order leads to inefficient characteristics.

### v3

@@ 8.2 Scene and Demonstration Generation
Simulation scenes can be created manually. However, this approach is time-consuming and labor-intensive. As a result, automated or semi-automated scene generation methods are more preferred (Deitke et al. (2022)). Two methods can be used. Real-to-Sim method converts real scenes to simulation. Automated generation method generates simulation scenes automatically without real-world observation. Real-to-Sim method can accurately mimic the real world, but it limits the diversity of scenes. The automated generation method can create more diverse scenes and increase the variety of collected demonstrations.
The Real-to-Sim method directly refers to a digital twin. The Real-to-Sim method utilizes 3D reconstruction technology (in Sec. 6) or inverse graphics (Chen et al. (2024c)) to create the real-world scene in a virtual environment (Torne et al. (2024)). But, 3D reconstruct scene is static environment where objects lack real-world physical properties, such as material, mass and friction coefficients, and are non-interactive (Torne et al. (2024)). The inverse graphics method, such as URDFormer (Chen et al. (2024c)), directly generates interactive simulation environment and articulated objects from input RGB image. Compared to 3D-generated methods, it reduces human involvement and produces interactive simulation environment. However, it lacks physical plausibility and fails to address the mismatch between the generated object’s physical properties and the real world. Previous research on physical parameter identification of real-world objects can be categorized into estimation from interaction (Seker and Kroemer (2024); Bohg et al. (2017); Xu et al. (2019)), estimation from demonstrations (Torne et al. (2024)). PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to estimate physical object attributes, such as material, from observation.
GRS (Zook et al. (2024)) employs SAM2 for object segmentation from RGBD image and utilizes VLMs to describe and match objects with simulation-ready assets. This approach combines the strengths of 3D reconstruction and inverse graphics methods. It ensures the credibility of 3D-generated methods and allowing objects in the scene to interact. However, it is impossible for the assert dataset to fully cover objects in the real world. ACDC (Dai et al. (2024)) defines a digital cousin concept. Unlike a digital twin, it does not directly replicate a real-world counterpart. However, it retains similar geometric and semantic features by using similar asserts when the assert dataset does not include real-world objects. Constructing an interact assert dataset often requires manual design by the creator or human-assisted interactive object generation.
The automated generation methods can be categorized into rule-based and learning-based approaches. For instance, ProcTHOR (Deitke et al. (2022)) introduces a procedural generation pipeline for interactive scenes using rule-based constraints and statistical priors. However, the generated scenes often rely on pre-defined priors, resulting in unrealistic outcomes that hinder agent learning (Khanna et al. (2024)). To address this, PHYSCENE (Yang et al. (2024)) incorporates physical collision avoidance, object layouts, interactivity, and reachability metrics into a diffusion model. This approach enhances the physical plausibility and interactivity of generated scenes. Due to the prior knowledge of foundation models, there are current efforts to use foundation models for scene construction. RoboGen (Wang et al. (2023d)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023b)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library. However, using foundation models to automate the generation of scene’s physical plausibility still relies on VLM for judgment.
To achieve scene diversity, RoboCasa (Nasiriany et al. (2024)) uses Madjourney to generate texture images for kitchen scenes. At the same time, many studies focus on the automatic generation of assets, such as zero-1-to-3 (Wang et al. (2023d)), Luma.ai (Nasiriany et al. (2024)), LLaMA-Mesh (Wang et al. (2024d)). The modeling of the interaction environment above primarily focuses on articulated object modeling. Articulated objects can be created manually by designers or generated using procedural (Jiang et al. (2022); Liu et al. (2023b); Zhang et al. (2023d)) or human-assisted interactive methods (Torne et al. (2024)) after 3D scanning. They can also be generated automatically through inverse graphics (Chen et al. (2024c)). However, current automated methods for generating articulated object assets are limited to objects with few rotational joints. Real2Code (Mandi et al. (2024)) fine-tunes a CodeLlama model to process visual observation descriptions and then outputs joint predictions. This enables Real2Code to reconstruct complex articulated objects with up to 10 parts.
@@
To collect demonstrations in simulations, different approaches can be used based on task complexity. For simple tasks, like a two-finger gripper picking up a cube, a hard-coding method (Wang et al. (2022b)) can be used. However, for more complex tasks, remote teleoperation (Chen et al. (2024a)) or skill library (Ha et al. (2023)) should be employed. Building skill library can be done using reinforcement learning or gradient optimization methods. RoboGen (Wang et al. (2023d)) shows that gradient-based trajectory optimization is better for fine-grained manipulation tasks with soft bodies, like shaping dough into a specific form. On the other hand, reinforcement learning and evolutionary strategies are more effective for contact-rich tasks and continuous interactions with other components in the scene.
@@ 8.3 Sim-to-Real Gap Solutions
The methods discussed above assume that the target domain remains unchanged. However, many physical parameters of the same robot can change significantly. Factors like temperature, humidity, positioning, and wear and tear over time can all affect these parameters. This makes it harder to bridge the sim-to-real gap. To address this issue, DORA (Zhang et al. (2024b)) uses an information bottleneck principle. It aims to maximize the mutual information between the dynamics encoding and environmental data. At the same time, it minimizes the mutual information between the dynamics encoding and the behavior policy actions. Transic (Jiang et al. (2024a)) proposes a data-driven approach that enables successful sim-to-real transfer using a human-in-the-loop framework.
@@ 8.4 Dataset Augmentation
Current data augmentation can be mainly divided into scene-level and object-level. Scene-level refers to changing the layout of objects in the scene. For example, MimicGen (Mandlekar et al. (2023)) and DexMimicGen (Jiang et al. (2024b)) change the positions and orientations of objects, while CACTI (Mandi et al. (2022)) adds new, artificial objects to the scene. Current foundation models for dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s texture. GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
Summary
@@
Figure 8: Foundation Models for Manipulation Data Generation. Current mainstream simulators include Pybullet, MuJoCo, CoppeliaSim, NVIDIA Omniverse, and Unity. Meanwhile, learning-based generative models used as simulators have shown potential. Simulation environment generation can be classified into Real-to-Sim and Automated Generation methods. In the Real-to-Sim methods, assuming the object’s position is known, the main challenge lies in constructing the object’s 3D mesh. This can be achieved through scanning technique or by using VGM to generate the 3D mesh directly from RGB image (Chen et al. (2024c)). Additionally, GRS (Zook et al. (2024)) utilizes VLM to extract 3D object meshes corresponding to real-world object from assert database based on RGB image. In the Automated Generation methods, LLM can output scene descriptions or scene code based on task instruction. When the output is a scene description, VGM generates the objects and arranges them according to the description. Meanwhile, the generated scene need to be evaluated by VLM. (Wang et al. (2023d)). When the output is scene code, it directly generates the corresponding scene (Wang et al. (2023b)). However, this requires substantial prior knowledge of scene code within the task library. There are three methods for generating demonstrations in a scene: Hard-code, Teleoperation, and Skill Library. When building skill library, gradient optimization is effective in training skill for deformable tasks and reinforcement learning works better for contact-rich tasks (Wang et al. (2023d)). Solutions for the Sim-to-Real gap include System Identification, Domain Randomization, and Transfer Learning. For data augmentation, VFM is used to segment images first, and then VGM renders the object’s texture on the masked image.
@@ 9.1 What is the framework for general manipulation?
Based on the general manipulation definition and robot learning development history, this paper proposes a framework for a general manipulation capability. Given that the scenarios are static, the framework is designed in a modular, sequential manner. To facilitate module migration, it is preferable for each module to be plug-and-play. Given the current reliance on human-in-the-loop mechanisms in autonomous driving and medical robotics to ensure safety, this framework aims for human-robot interaction through corrective instruction to ensure the safety of manipulation actions. The corrective action can be collected into the dataset and then improve old skills through offline training.
9.1.3 The proposed framework limitations.
(1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates, precision and safety for specific tasks (Pang et al. (2023)). Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework proposed in this paper is based on the development of learning-based methods and the definition of general manipulation. The framework of brain-like cognitive research should also be explored.
@@
During robot execution, continuous human supervision is not always feasible. Hence, integrating real-time monitoring through parallel surveillance videos during robot execution could enhance safety. The framework in this paper does not explicitly denote this parallel safety monitoring module, as it resembles the post-conditions detection module. The post-conditions detection module analyzes the robot’s execution video to identify reasons for task failure, facilitating post-hoc correction to ensure task success. If the algorithm’s task execution safety is 80%, and the monitoring module predicts safety at 80% as well, the probability of risky movements reduces to 4%. Of course, for household robots, ensuring an over 99% safety rate is imperative. Initially, cloud-based monitoring of multiple robots by a single operator, with human intervention to correct erroneous behaviors, appears to be the best approach. This strategy not only reduces labor requirements but also ensures safety. Later, by gathering extensive data to improve model accuracy.